# FairPO: Robust Preference Optimization for Fair Multi-Label Learning

### v3

@@ Abstract
Multi-label classification (MLC) often exhibits performance disparities, especially for infrequent or sensitive label categories. We propose FairPO, a novel framework that integrates preference-based loss formulations with group-robust optimization to improve fairness in multi-label classification (MLC), particularly targeting underperforming and sensitive label groups. FairPO partitions labels into a privileged set, targeted for enhanced performance, and a non-privileged set, where baseline performance is maintained. For privileged labels, a preference-based loss, inspired by Direct Preference Optimization (DPO), encourages model scores for true positives to be significantly higher than for their confusing negative counterparts, and scores for true negatives to be significantly lower than for their confusing positive counterparts—addressing hard examples that challenge standard classifiers. For non-privileged labels, a constrained objective ensures performance does not degrade substantially below a reference model. These group-specific objectives are balanced using a Group Robust Preference Optimization (GRPO) formulation, adaptively mitigating bias. Our experiments also include FairPO variants built on recent reference-free preference optimization techniques, namely Contrastive Preference Optimization (CPO) and Simple Preference Optimization (SimPO), which further highlight FairPO’s versatility [^1] Our code is available at GitHub: https://anonymous.4open.science/r/FairPO.
@@ 1 Introduction
Multi-label classification (MLC)—assigning a subset of labels from a universe $\mathcal{T}$ to an instance xx—is a pervasive task in fields like image annotation (Wang et al., 2016) and document categorization (Zangari et al., 2024; Schietgat et al., 2010). Typically, models are trained by minimizing per-label binary cross-entropies (BCE) over a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ (Zhang and Zhou, 2014; Sorower, 2010; Tarekegn et al., 2024). However, this global optimization often creates fairness issues (Mehrabi et al., 2022a; Chouldechova, 2017), as imbalanced label frequencies and varying importance lead to disparate performance (Mehrabi et al., 2022b). For instance, a model might excel on common labels while consistently failing on rare but critical ones. This disparity is particularly problematic when certain labels correspond to protected attributes or groups, demanding equitable treatment.
Figure 1: FairPO’s methodology
A key challenge lies in the model’s ability to discriminate effectively, especially for hard examples. Standard losses like BCE (Ruby and Yendapalli, 2020) may provide insufficient signal for nuanced distinctions. For a given label $l$, if it is truly positive ($y_{il}=+1$), the model might still assign a low score $m(x_{i};\mathbf{w}_{l})$ or, more critically, assign a higher score to an incorrect label $k$ such that $y_{ik}=0$, i.e., a confusing negative that the model mistakenly ranks above the true label. Conversely, if label $l$ is truly negative ($y_{il}=0$), the model might erroneously assign it a high score $m(x_{i};\mathbf{w}_{l})$ (a confusing positive). Existing fairness interventions often target single-label settings (Zafar et al., 2017; Hardt et al., 2016; Dwork et al., 2011) and are not readily adapted to the multi-label context and these symmetric discriminative challenges.
To address this, we introduce FairPO (Fair Preference Optimization), a novel framework that integrates preference-based learning with group robust optimization. Inspired by the success of Direct Preference Optimization (DPO) in aligning models with human preferences (Rafailov et al., 2024), FairPO recasts parts of the MLC task as learning explicit preferences. We partition the label set $\mathcal{T}$ into a privileged set $\mathcal{P}$, where enhanced, fair performance is paramount, and a non-privileged set $\bar{\mathcal{P}}=\mathcal{T}\setminus\mathcal{P}$, where the goal is to maintain robust baseline performance.
For all the privileged labels $l\in\mathcal{P}$, FairPO employs a conditional objective. If a confusing set exists for $(x_{i},l)$—meaning there are confusing negatives $k\in S_{il}^{\text{neg}}$ (where $y_{ik}=0,m(x_{i};\mathbf{w}_{k})\geq m(x_{i};\mathbf{w}_{l})$ when $y_{il}=+1$) or confusing positives $k^{\prime}\in S_{il}^{\text{pos}}$ (where $y_{ik^{\prime}}=+1,m(x_{i};\mathbf{w}_{k^{\prime}})\leq m(x_{i};\mathbf{w}_{l})$ when $y_{il}=0$)—a preference loss (inspired by DPO (Rafailov et al., 2024), SimPO (Meng et al., 2024), or CPO (Xu et al., 2024)) is applied. This loss encourages $m(x_{i};\mathbf{w}_{l})\gg m(x_{i};\mathbf{w}_{k})$ for the $(l,k)$ pair if $y_{il}=+1$, or $m(x_{i};\mathbf{w}_{l})\ll m(x_{i};\mathbf{w}_{k^{\prime}})$ for the $(l,k^{\prime})$ pair if $y_{il}=0$. If no such confusing examples are found for $(x_{i},l)$, a standard BCE loss for label $l$ is applied to ensure consistent learning. For non-privileged labels $j\in\bar{\mathcal{P}}$, FairPO enforces a constraint that their classification loss $\ell(\mathbf{w}_{j};x_{i},y_{ij})$ does not substantially exceed that of a reference model $\hat{\mathbf{w}}_{j}$, i.e., $\ell(\mathbf{w}_{j})\leq\ell(\hat{\mathbf{w}}_{j})+\epsilon$. To manage the trade off between these two distinct objectives, FairPO leverages a group specific robust optimization technique—Group Robust Preference Optimization (GRPO) (Ramesh et al., 2024). The overall learning problem is formulated as a standard minimax optimization: $\min_{\{\mathbf{w}_{t}\}_{t\in\mathcal{T}}}\max_{\alpha\in\Delta_{1}}\left[\alpha_{\mathcal{P}}\mathcal{L}_{\mathcal{P}}+\alpha_{\bar{\mathcal{P}}}\mathcal{L}_{\bar{\mathcal{P}}}\right]$, where $\mathcal{L}_{\mathcal{P}}$ and $\mathcal{L}_{\bar{\mathcal{P}}}$ are the aggregate losses for the privileged and non-privileged groups respectively, and $\alpha=(\alpha_{\mathcal{P}},\alpha_{\bar{\mathcal{P}}})$ are adaptive weights in the simplex $\Delta_{1}$. This dynamically balances training, preventing performance degradation in one group for gains in another, thereby promoting fairness across the labels (see Figure 1).
Our primary contributions are a novel framework for fair MLC using preference signals, a conditional objective targeting hard examples, and a robust optimization strategy to manage fairness trade-offs. Our comprehensive experiments show that FairPO, particularly the CPO-variant, achieves significant gains—for instance, up to 3.44% mAP on the least frequent labels of COCO (Lin et al., 2015)—while robustly maintaining performance on non-privileged labels. We also outline the framework’s applicability to attribute generation in Appendix A.
2 Preliminaries: Preference Optimization Methods
Direct Preference Optimization (DPO): DPO (Rafailov et al., 2024) directly optimizes a policy $\pi_{\theta}$ using preference pairs $(x,y_{w},y_{l})$, where $y_{w}$ is preferred over $y_{l}$ for prompt $x$. Assuming a Bradley-Terry preference model tied to an implicit reward function related to $\pi_{\theta}$ and a reference policy $\pi_{\text{ref}}$, DPO maximizes the likelihood of observed preferences, resulting in the loss:
Group Robust Preference Optimization (GRPO): GRPO (Ramesh et al., 2024) extends preference optimization to handle diverse preference groups $\{D_{g}\}_{g=1}^{K}$. Instead of minimizing the average loss, GRPO minimizes the worst-case loss across groups using a robust objective:
Simple Preference Optimization (SimPO): SimPO (Meng et al., 2024) aims to align the implicit reward with generation metrics and eliminates the need for $\pi_{\text{ref}}$. It uses the length-normalized average log-likelihood as the reward: $r_{\text{SimPO}}(x,y)=\frac{\beta}{|y|}\log\pi_{\theta}(y|x)$. It also introduces a target margin $\gamma>0$ into the preference model. The resulting SimPO loss is:
Contrastive Preference Optimization (CPO): CPO (Xu et al., 2024) also removes the dependency on $\pi_{\text{ref}}$ for efficiency, approximating the DPO objective. It combines a reference-free preference loss with a negative log-likelihood (NLL) regularizer on preferred responses $y_{w}$ to maintain generation quality:
$L_{\text{CPO}}(\pi_{\theta})$ $=L_{\text{prefer}}+L_{\text{NLL}}.$ (7)
3 Methodology: Fair Preference Optimization (FairPO)
@@ 3.1 Problem Setup and Fairness Goals
Given a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, we fine-tune parameters $\mathbf{w}_{t}$ for a per-label classifier that assigns a score $m(x_{i};\mathbf{w}_{t})$ for each label $t\in\mathcal{T}$. Our framework leverages a reference model with parameters $\hat{\mathbf{w}}_{t}$, typically obtained from standard supervised fine-tuning (SFT).
Our framework addresses the problem of group fairness, where the goal is to ensure equitable performance across different predefined groups. In our MLC context, we define these groups not by instance attributes (e.g., demographics), but by partitioning the labels themselves into a privileged set $\mathcal{P}$ (e.g., rare, critical, or historically underperforming labels) and a non-privileged set $\bar{\mathcal{P}}$. Our fairness objective is a nuanced form of equitable treatment: we aim to significantly improve performance for the underperforming privileged group while ensuring that performance for the non-privileged group is not substantially harmed. This goal of targeted improvement without undue harm to other groups is enforced by our robust optimization framework.
For any privileged label $l\in\mathcal{P}$, the model must accurately discriminate its true state from confusing alternatives. For any non-privileged label $j\in\bar{\mathcal{P}}$, performance should not degrade significantly from the reference model. To formalize this, we define two types of confusing sets for a privileged label $l\in\mathcal{P}$: the set of confusing negatives, $S_{il}^{\text{neg}}=\{k\in\mathcal{T}\mid y_{ik}=0\text{ and }m(x_{i};\mathbf{w}_{k})\geq m(x_{i};\mathbf{w}_{l})\}$ when $y_{il}=+1$, and the set of confusing positives, $S_{il}^{\text{pos}}=\{k\in\mathcal{T}\mid y_{ik}=+1\text{ and }m(x_{i};\mathbf{w}_{k})\leq m(x_{i};\mathbf{w}_{l})\}$ when $y_{il}=0$. The overall confusing set is $S_{il}=S_{il}^{\text{neg}}\cup S_{il}^{\text{pos}}$.
3.2 Objective for Privileged Labels ($l\in\mathcal{P}$)
Confusing Examples Exist ($S_{il}\neq\emptyset$): We employ a DPO-inspired (Rafailov et al., 2024) preference loss. If $y_{il}=+1$ and $k\in S_{il}^{\text{neg}}$ is sampled, we prefer $l$ over $k$ ($l\succ k$). The DPO term is:
$h_{\mathbf{w}}(x_{i},l,k)=\left(\log\frac{m(x_{i};\mathbf{w}_{l})}{m(x_{i};\hat{\mathbf{w}}_{l})}\right)-\left(\log\frac{m(x_{i};\mathbf{w}_{k})}{m(x_{i};\hat{\mathbf{w}}_{k})}\right).$ (8)
The loss contribution is $\log\sigma(\beta\cdot h_{\mathbf{w}}(x_{i},l,k))$.
If $y_{il}=0$ and $k\in S_{il}^{\text{pos}}$ is sampled, we prefer $l$ over $k$ (i.e., the true negative $l$ is preferred over the confusing positive $k$, meaning $m(x_{i};\mathbf{w}_{l})$ should be lower than $m(x_{i};\mathbf{w}_{k})$ relative to the reference). The DPO term is (note the swapped order for $k$ and $l$ to reflect $k\succ l$ effectively aiming for $m(x_{i};\mathbf{w}_{l})<m(x_{i};\mathbf{w}_{k})$):
$h_{\mathbf{w}}(x_{i},k,l)=\left(\log\frac{m(x_{i};\mathbf{w}_{k})}{m(x_{i};\hat{\mathbf{w}}_{k})}\right)-\left(\log\frac{m(x_{i};\mathbf{w}_{l})}{m(x_{i};\hat{\mathbf{w}}_{l})}\right).$ (9)
The loss contribution is $\log\sigma(\beta\cdot h_{\mathbf{w}}(x_{i},k,l))$.
No Confusing Examples ($S_{il}=\emptyset$): We revert to a standard base classification loss (e.g., Binary Cross-Entropy, BCE) for label $l$. This BCE fallback ensures that the model continues to receive learning signals even for labels without any identified confusing counterparts, thereby avoiding training stagnation.
$\ell_{\text{BCE}}(x_{i},y_{il};\mathbf{w}_{l})=-[y_{il}\log m(x_{i};\mathbf{w}_{l})+(1-y_{il})\log(1-m(x_{i};\mathbf{w}_{l}))].$ (10)
The overall Privileged Loss $\mathcal{L}_{\mathcal{P}}$ is the expectation over these conditional losses:
$\mathcal{L}_{\mathcal{P}}(\{\mathbf{w}_{t}|t\in\mathcal{T}\},\{\hat{\mathbf{w}}_{t}|t\in\mathcal{T}\})$ $=\mathbb{E}_{(x_{i},l)\text{ s.t. }l\in\mathcal{P}}\left[\mathbf{1}_{S_{il}\neq\emptyset}\cdot(-\log\sigma(\beta\cdot h_{\text{DPO}}))\right$.
$\quad\left.+\mathbf{1}_{S_{il}=\emptyset}\cdot\ell_{\text{BCE}}(x_{i},y_{il};\mathbf{w}_{l})\right]$ (11)
where $h_{\text{DPO}}$ refers to the appropriate term from Eq. 8 or Eq. 9. The hyperparameter $\beta$ controls the preference strength.
SimPO-Inspired Loss: We also adapt insights from Simple Preference Optimization (SimPO) (Meng et al., 2024). Original SimPO for generative models uses length-normalized log-likelihoods and a target margin $\gamma$ in its preference scoring. Since our multi-label classification (MLC) setup deals with individual label scores $m(x_{i};\mathbf{w}_{t})$ where sequence length is inapplicable, we adapt SimPO’s core concept of a target margin $\gamma$ to our context. This results in a reference-free preference component that aims to ensure a minimum separation for preferred label scores, omitting length normalization but retaining the margin:
$L_{\text{pref}}^{\text{SimPO}}(x_{i},l,k)=-\log\sigma\left(\beta\left(\log\frac{m(x_{i};\text{preferred})}{m(x_{i};\text{dispreferred})}\right)-\gamma\right),$ (12)
where $(m(x_{i};\text{preferred}),m(x_{i};\text{dispreferred}))$ are $(m(x_{i};\mathbf{w}_{l}),m(x_{i};\mathbf{w}_{k}))$ if $y_{il}=+1$ and $k\in S_{il}^{\text{neg}}$, or $(m(x_{i};\mathbf{w}_{k}),m(x_{i};\mathbf{w}_{l}))$ if $y_{il}=0$ and $k\in S_{il}^{\text{pos}}$. The term $\beta\log(\cdot)$ captures the relative preference, while $\gamma$ enforces a desired separation margin. This $L_{\text{pref}}^{\text{SimPO}}$ replaces the DPO term in Eq. 11 when $S_{il}\neq\emptyset$, with the BCE fallback (Eq. 10) for $S_{il}=\emptyset$ remaining. More broadly, this adaptation leverages SimPO’s margin mechanism for more distinct score separation in challenging MLC cases.
CPO-Inspired Loss: Adapting Contrastive Preference Optimization (Xu et al., 2024). If $S_{il}\neq\emptyset$, the preference component is:
$L_{\text{pref}}^{\text{CPO}}(x_{i},l,k)=-\log\sigma\left(\beta\left(\log\frac{m(x_{i};\text{preferred})}{m(x_{i};\text{dispreferred})}\right)\right).$ (13)
(using the same preferred/dispreferred logic as SimPO). CPO also includes an NLL regularizer. Thus, the overall CPO-inspired privileged loss, $\mathcal{L}_{\mathcal{P}}^{\text{CPO}}$, could be formulated by combining this preference term (when $S_{il}\neq\emptyset$) with a consistent NLL/BCE component for all true ground truth labels within $\mathcal{P}$:
[…]
where $k_{s}$ is a confusing example sampled from $S_{il}$, and $\lambda_{\text{CPO}}$ is a weighting hyperparameter. Note that the $\ell_{\text{BCE}}$ term is always applied, serving as both the NLL regularizer and the fallback for $S_{il}=\emptyset$ if the first term becomes zero. These alternatives offer different ways to model preferences. The non-privileged loss $\mathcal{L}_{\bar{\mathcal{P}}}$ (Eq. 17) remains unchanged regardless of the privileged loss choice.
3.3 Constrained Objective for Non-Privileged Labels ($j\in\bar{\mathcal{P}}$)
For non-privileged labels $j\in\bar{\mathcal{P}}$, our goal is to maintain performance relative to the reference model $\hat{\mathbf{w}}_{j}$, preventing degradation due to the focus on privileged labels. We use a standard base classification loss $\ell(\mathbf{w}_{j};x_{i},y_{ij})$ (e.g., Binary Cross-Entropy) defined as:
$\ell(\mathbf{w}_{j};x_{i},y_{ij})$ $=-[y_{ij}\log m(x_{i};\mathbf{w}_{j})+(1-y_{ij})\log(1-m(x_{i};\mathbf{w}_{j}))],$ (15)
$\ell(\hat{\mathbf{w}}_{j};x_{i},y_{ij})$ $=-[y_{ij}\log m(x_{i};\hat{\mathbf{w}}_{j})+(1-y_{ij})\log(1-m(x_{i};\hat{\mathbf{w}}_{j}))].$ (16)
The Non-Privileged Loss $\mathcal{L}_{\bar{\mathcal{P}}}$ employs a hinge mechanism, penalizing the model only if its loss $\ell(\mathbf{w}_{j})$ exceeds the reference model’s loss $\ell(\hat{\mathbf{w}}_{j})$ by more than a predefined slack margin $\epsilon\geq 0$:
$\mathcal{L}_{\bar{\mathcal{P}}}(\{\mathbf{w}_{t},\hat{\mathbf{w}_{t}}|t\in\mathcal{T}\})=\mathbb{E}_{(x_{i},j)\text{ s.t. }j\in\bar{\mathcal{P}}}\left[\max(0,\ell(\mathbf{w}_{j};x_{i},y_{ij})-\ell(\hat{\mathbf{w}}_{j};x_{i},y_{ij})-\epsilon)\right].$ (17)
This ensures $\mathbf{w}_{j}$ is primarily updated only if performance on label $j$ drops significantly below the reference baseline plus $\epsilon$.
3.4 Group Robust Optimization Formulation
To effectively balance the objective for privileged labels ($\mathcal{L}_{\mathcal{P}}$) and the constraint for non-privileged labels ($\mathcal{L}_{\bar{\mathcal{P}}}$), we employ the Group Robust Preference Optimization (GRPO) framework (Ramesh et al., 2024). We define two distinct groups corresponding to our label partitions:
  • $G_{\mathcal{P}}$: Associated with the Privileged Loss $\mathcal{L}_{\mathcal{P}}$, focusing on triplets $(x_{i},l,k)$.
  • $G_{\bar{\mathcal{P}}}$: Associated with the Non-Privileged Loss $\mathcal{L}_{\bar{\mathcal{P}}}$, focusing on pairs $(x_{i},j)$.
The FairPO objective becomes the GRPO minimax problem:
$\min_{\{\mathbf{w}_{t}|t\in\mathcal{T}\}}\max_{\alpha_{\mathcal{P}},\alpha_{\bar{\mathcal{P}}}\,\geq\,0,\,\alpha_{\mathcal{P}}+\alpha_{\bar{\mathcal{P}}}\,=\,1}\left[\alpha_{\mathcal{P}}\mathcal{L}_{\mathcal{P}}(\{\mathbf{w}_{t},\hat{\mathbf{w}_{t}}|t\in\mathcal{T}\})+\alpha_{\bar{\mathcal{P}}}\mathcal{L}_{\bar{\mathcal{P}}}(\{\mathbf{w}_{t},\hat{\mathbf{w}_{t}}|t\in\mathcal{T}\})\right].$ (18)
Here, $\alpha_{\mathcal{P}}$ and $\alpha_{\bar{\mathcal{P}}}$ are adaptive weights representing the importance assigned to each group’s loss. The inner maximization finds the worst-case distribution over group losses (by increasing the weight $\alpha$ for the group with higher current loss), while the outer minimization seeks model parameters $\{\mathbf{w}_{t}|t\in\mathcal{T}\}$ that perform well even under this worst-case weighting. This formulation inherently promotes robustness and fairness by preventing the optimization from disproportionately favouring one group at the expense of the other.

### v4

@@ Abstract
Multi-label classification (MLC) often suffers from performance disparities across labels. We propose FairPO, a framework combining preference-based loss and group-robust optimization to improve fairness by targeting underperforming labels. FairPO partitions labels into a privileged set for targeted improvement and a non-privileged set to maintain baseline performance. For privileged labels, a DPO-inspired preference loss addresses hard examples by correcting ranking errors between true labels and their confusing counterparts. A constrained objective maintains performance for non-privileged labels, while a Group Robust Preference Optimization (GRPO) formulation adaptively balances both objectives to mitigate bias. We also demonstrate FairPO’s versatility with reference-free variants using Contrastive (CPO) and Simple (SimPO) Preference Optimization [^1] Our code is available at GitHub: https://github.com/soumenkm/FairPO.
00footnotetext: Equal contribution
@@ 1 Introduction
The Challenge of Performance Disparity in MLC.
Standard multi-label classification (MLC) models assume all labels are equal, minimizing an aggregate loss over the entire label set (kiritchenko2006hierarchical; DBLP:journals/corr/WeiXLZ0G16; schietgat2010predicting). This assumption fails in real-world scenarios where data asymmetries cause significant performance disparities. These disparities stem from several factors, including class imbalance, where models favor frequent labels over rare ones (Charte2015AddressingII; 6471714), and varying real-world importance, such as distinguishing a critical disease from a benign one (rajpurkar2017chexnetradiologistlevelpneumoniadetection; 10368872). Disparities also arise from semantic complexity, where models learn simple concepts but fail on nuanced ones. Consequently, training becomes dominated by the easy majority, yielding systems that are unreliable for the very cases that matter most (swayamdipta2020datasetcartographymappingdiagnosing; pleiss2020identifyingmislabeleddatausing).
Limitations of Conventional Methods.
Conventional methods are limited by the aggregate nature of the Binary Cross-Entropy (BCE) loss, which allows the learning signal for underrepresented labels to be drowned out by easy ones (bce_loss; Charte2015AddressingII). While solutions like Focal Loss help, they are fundamentally restricted because they treat each label’s prediction in isolation (lin2018focallossdenseobject). They do not teach the model to perform discrimination by resolving confusion between a true label and a specific negative alternative. This reveals a gap for a framework that can learn relative preferences, improving performance on difficult labels while maintaining it for others. A preference-based objective forces the model to learn that a true label’s score is explicitly preferred over that of a confusing competitor.
Our Approach: The FairPO Framework.
We introduce FairPO (Fair Preference Optimization), a framework to manage performance disparities by partitioning labels into a privileged group for targeted improvement and a non-privileged group for performance maintenance. We prioritize learning on the privileged group using a preference-based objective, a novel adaptation of techniques from generative model alignment for a discriminative task. For each privileged true label, FairPO dynamically identifies its confusing counterparts, incorrect labels with misleadingly high scores and trains the model to prefer the true label, creating a large discriminative margin. Concurrently, a constrained objective safeguards the non-privileged group against performance degradation relative to a reference model. These competing objectives are dynamically balanced using a Group Robust Preference Optimization (GRPO) formulation (ramesh2024group).
Figure 1: FairPO Framework
Application and Contributions.
Our general FairPO framework can address any MLC task requiring label prioritization due to factors like class imbalance or real-world importance. To demonstrate its effectiveness, we apply it to class imbalance, defining infrequent tail labels as the privileged group and frequent head labels as the non-privileged one. Our main contributions are:
  1. 1.
    We propose FairPO, a novel framework for targeted performance improvement in MLC applicable to various sources of disparity.
  2. 2.
    We are the first to adapt preference optimization techniques (meng2024simposimplepreferenceoptimization; xu2024contrastivepreferenceoptimizationpushing) from generative modeling to create discriminative objectives for MLC.
  3. 3.
    We successfully apply GRPO to solve the trade-off between objectives for the privileged and non-privileged groups, a novel application in this context.
Experiments show FairPO significantly boosts performance on infrequent labels by up to 3.59% mAP while robustly maintaining performance on frequent ones, outperforming several baselines.
2 Methodology: The FairPO Framework
Our methodology builds upon a foundation of preference optimization techniques. To provide the necessary background on these methods, namely DPO, CPO, SimPO, and GRPO, we have moved the detailed review of their core formulations to Appendix A.
@@ 2.1 Problem Setup and Fairness Goals
Let us consider a standard multi-label classification setting with a dataset $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$, where $\mathbf{x}_{i}$ is an input instance and $y_{i}\in\{0,1\}^{|\mathcal{T}|}$ is the corresponding multi-label vector over a universe of $|\mathcal{T}|$ labels. The goal is to train a model, parameterized by weights $\mathbf{w}$, that produces a set of per-label scores $m(\mathbf{x}_{i};\mathbf{w}_{t})$ for each label $t\in\mathcal{T}$. Our framework also utilizes a pre-trained reference model with parameters $\hat{\mathbf{w}}_{t}$.
Conventional methods often train such models by minimizing a single, aggregate loss where each label contributes equally. However, as motivated in our introduction, real-world applications frequently demand a more nuanced approach where certain labels are prioritized. Our framework, FairPO, is explicitly designed for these scenarios. The core idea is to partition the total label set $\mathcal{T}$ into two disjoint subsets:
  • A privileged group $\mathcal{P}\subset\mathcal{T}$, which contains labels for which we seek to significantly enhance the model’s performance.
  • A non-privileged group $\bar{\mathcal{P}}=\mathcal{T}\setminus\mathcal{P}$, which contains the remaining labels, for which our objective is to robustly maintain at least a baseline level of performance.
This partitioning is a key strength of our framework’s general design. The criteria for assigning labels to the privileged group are flexible and can be adapted to the specific problem at hand, such as using label frequency for class imbalance, domain knowledge for real-world importance, or annotation consistency for data quality issues. In this work, to demonstrate FairPO’s efficacy, we focus on the class imbalance problem, where the privileged group consists of the least frequent labels. We defer the precise details of this setup to Section 3.
2.2 FairPO Objectives
FairPO’s methodology is built on three core components: a conditional objective for privileged labels, a constrained objective for non-privileged labels, and a robust optimization framework to balance them. The detailed mathematical formulations for each component are provided in Appendix B.
Objective for Privileged Labels ($l\in\mathcal{P}$): To improve performance on critical labels, we employ a conditional objective designed to target hard-to-discriminate cases. When a confusing set of labels exists for an instance (i.e., high-scoring negatives or low-scoring positives), a preference loss inspired by DPO (rafailov2023direct) is applied to enforce correct relative rankings. In the absence of such confusing examples, the objective reverts to a standard BCE loss, which acts as a crucial fallback for stable training. Our framework is versatile and also supports reference-free preference loss variants inspired by CPO (xu2024contrastivepreferenceoptimizationpushing) and SimPO (meng2024simposimplepreferenceoptimization).
Objective for Non-Privileged Labels ($j\in\bar{\mathcal{P}}$): To maintain baseline performance on the remaining labels, we use a constrained objective. This objective employs a hinge mechanism that only incurs a penalty if the model’s performance on a label drops significantly below that of a reference model by a predefined slack margin $\epsilon$. This acts as a protective measure, preventing substantial performance degradation for the non-privileged group.
Group Robust Optimization: These two distinct objectives for the privileged and non-privileged groups are adaptively balanced using the Group Robust Preference Optimization (GRPO) framework (ramesh2024group). The overall learning problem is formulated as the following minimax objective:
$\min_{\{\mathbf{w}_{t}|t\in\mathcal{T}\}}\max_{\alpha_{\mathcal{P}}+\alpha_{\bar{\mathcal{P}}}=1}\left[\alpha_{\mathcal{P}}\mathcal{L}_{\mathcal{P}}(\cdot)+\alpha_{\bar{\mathcal{P}}}\mathcal{L}_{\bar{\mathcal{P}}}(\cdot)\right].$ (1)
This formulation dynamically adjusts the training focus between the privileged and non-privileged groups, seeking a solution that is robust to the worst-case group loss and thereby managing the fairness-performance trade-off.